# Impact of Charge Transfer Inefficiency on transit light-curves

### v1

@@ 2.5 Bias on transit depth measurements
To assess the impact of CTI on the transit method, we simulated representative light-curves for both Earth- and Jupiter-sized planets transiting a solar-type star. The intrinsic stellar transit signal is generated using the PLATO Solar-like Light-curve Simulator [^2] https://sites.lesia.obspm.fr/psls/ (Samadi et al., 2019), which is based on the analytical model by Mandel & Agol (2002) and its Python implementation by Ian Crossfield (MIT) [^3] http://www.astro.ucla.edu/~ianc/.
The flux variations over time are interpolated onto a grid of 2000 time steps. The stellar image is modelled using a Gaussian Point Spread Function (PSF) with a characteristic width of $0.6$ pixel. The total flux is scaled to match the expected value for a given stellar magnitude, following the flux–magnitude relation from Marchiori et al. (2019). Finally, this flux distribution is modulated by the intrinsic transit light-curve.
We added a constant background contribution from zodiacal light of $45\penalty 10000\ \mathrm{e^{-}\,s^{-1}\,pixel^{-1}}$ (Marchiori et al., 2019). Given the integration time of $21$s for the PLATO CCD, the net background value is $\approx 1000\penalty 10000\ \mathrm{e^{-}\,pixel^{-1}}$. A higher background fills traps that would otherwise capture flux from the target star, thereby reducing the CTI effect. The aperture size is determined using our optimal aperture mask routine, following the approach of Marchiori et al. (2019). Finally from the simulated light-curve, a transit depth is extracted.
@@
We calculated the bias on transit depth due to CTI for cases corresponding to nominal mission lifetime of 4 years and extended mission lifetime of 8 years. We also incorporated positional dependence, based on the imagette’s location in a region of high or low CTI. We evaluated three representative CCD positions, bottom left (A), top left (B), and top right (C), corresponding to different trap density regimes derived from Eq.7 (see Fig.4). The bottom left corner (A) is the region with the lowest cumulative trap density. Conversely, the top right corner (C) has the highest cumulative trap density. The top left corner (B) represents an intermediate case.
@@
To quantify the impact of CTI on transit light-curves, we defined the photometric bias, $\Delta$, as the difference between the pristine transit depth (no CTI) and the CTI-affected depth, expressed in both pixels-per-million (ppm) and percentage. Table 2 reports the impact of CTI on the transit depth of an Earth-sized planet transiting a solar-type star with $m_{V}=11$ under different mission durations and trap density scenarios. For completeness, it also presents the corresponding bias for a bright star ($m_{V}=9$) and a faint star ($m_{V}=13$) in a region of high CTI (position C) during the extended 8-year mission. It is worth noting that these calculations were done for the worst CCD (among all the cameras), i.e. for the CCD with the highest maximum in the TNID.
Table 2: CTI depth across mission durations, detector positions, and magnitudes. The pristine transit depth is equal to $105.8$ ppm (typical value for an Earth-sized planet orbiting a solar-type star).
@@
The origin of this bias is the CTI-induced charge loss, which does not scale linearly with the stellar flux. As a result, the deferred charge release affects the in-transit and out-of-transit measurements differently, thereby influencing the measured transit depth. This non-linear behaviour is the root cause of the photometric bias seen in the observed light-curve.
We found that the horizontal drift has no impact on the bias, as the light-curves have been de-trended beforehand. This will also be the case for the PLATO light-curves (however, see the discussion in Sect. 5.2).
Figure 5: Comparison of transit light-curves for the worst-case scenario: an 8-year mission at high trap density (position C) and $m_{V}=11$. The solid orange curve includes CTI effects, while the dashed blue curve shows the CTI-free case. CTI-induced charge loss introduces a photometric bias of $3.95\%$.
PLATO science requirements dictate that the planet-to-star radius ratio $R_{\mathrm{p}}/R_{\mathrm{s}}$ must be measured with a relative precision better than 2 % for an Earth-sized planet transiting a G-type star (goal: $m_{V}=11$) (PLATO Science Working Team, 2019). Because the transit depth $\delta$ is proportional to the square of this ratio, standard error propagation yields
@@ 3 CTI calibration strategy
This section outlines the calibration strategy employed in this study, which relies on full or partial images acquired at the beginning of each quarter (i.e., every three months). First, we define an empirical model to determine the number of trap species and to derive initial estimates for their release time constants $\tau_{\mathrm{r,k}}$. Next, we address and remove the contribution from smearing. Finally, we adopt a polynomial model to describe the spatial variation in the trap density, reproducing the radial dependence discussed in Sect. 2.4. We then apply the EPER method (Prod’homme et al., 2016) to constrain the polynomial coefficients of the trap density model $a_{\mathrm{p,k}}$, along with the $\beta$ parameter and the release time constants $\tau_{\mathrm{r,k}}$. The overall procedure is summarised in Fig. 6.
Figure 6: Workflow of the CTI calibration strategy: initial estimates for the release time constants $\tau_{\mathrm{r,k}}$ are obtained by fitting an empirical model to the overscan signal; smearing is corrected via a column-wise exponential-plus-constant fit; the mean trap density is modelled as a quadratic polynomial to reproduce its radial dependence; and the EPER method is applied, in conjunction with a CTI correction algorithm, to iteratively fit the polynomial coefficients, the well-fill power-law index $\beta$, and refine the release time constants $\tau_{\mathrm{r,k}}$.
@@ 3.1 Calibration dataset
As PLATO is not yet in-flight, we used a simulated CCD image containing a stellar field for the calibration procedure. This image is generated with PLATOSim (Jannsen et al., 2024), a simulation tool designed to produce photometric time series and CCD images tailored to the PLATO mission. The simulated image includes several sources of noise—smearing, photon noise, and readout noise—to approximate realistic observational conditions. To improve the signal-to-noise ratio (S/N), we averaged over 10 exposures. The upper limit of 10 exposures is decided by the in-flight operational constraints.
@@ 3.2 Overscan analysis and CTI signature
Parallel overscan rows are virtual rows beyond the CCD’s physical imaging area, accessed after the last physical row is read out. Hereafter, “overscan” refers to this parallel overscan region. These rows contain deferred charges that were not recorded during readout because of trap capture, providing a diagnostic window to constrain the CTI signature. Fig. 7 shows the overscan signal, summed across all the columns for a given row index. Its shape reflects the delayed release of charge by different trap species: those with short release times dominate near the start, whereas slower species extend further into the overscan region. The slowest traps are of primary concern, as their release occurs beyond the photometric aperture.
Figure 7: Sum of signal in each overscan row across all columns as a function of row index. The profile follows an exponential decay, with physical rows located before overscan row index 0.
@@ 3.3 Initial estimates for the trap release time (Step 1)
From the profile of the signal in Fig. 7, a multi-exponential function is used to (empirically) model the signal contained in the overscan region. It is illustrated in the Eq 9 as
@@
where the variables are defined as follows:
  • $A_{k}$: Amplitude for trap species $k$, corresponding to the range (max - min) of the signal.
  • $b_{k}$: Decay rate for species $k$, inversely related to the release time (see Eq 10),
  • $C$: Constant offset (sum of smearing in each individual column),
  • $i$: Row index along the CCD column (parallel transfer direction),
  • $n$: Total number of trap species,
  • $Y_{\text{CTI}}(i)$: Signal level (in e-/pix) at row $i$, including CTI trailing effects.
The relationship between the decay rate $b_{k}$ and release time ${\tau}_{r,k}$ is given by
${\tau}_{r,k}=\frac{1}{b_{k}}\cdot t_{\text{transfer\ cadence}}.$ (10)
We designed an iterative fitting strategy that progressively increases the number of species until the solution converges within a predefined threshold. Applying this strategy to our simulated image yielded a maximum of four trap species, with a final reduced $\chi^{2}$ of $1.141$. The first estimates for the release time values and decay constants for these species, ordered from fastest to slowest, are listed in Table 4. We provide the associated uncertainties for the release times and classify each species according to its characteristic release timescale: fast, intermediate, or slow. The corresponding fit results, along with the residuals, are presented in Fig. 8. The difference between our best-fit release time values (Table 4) and the reference values from Prod’homme et al. (2016) (Table 1) arises from the distinct methodologies employed in each case. Our simulations include a realistic stellar field, which influences the empirical model used to derive best-fit release times, unlike the flat-field image adopted by Prod’homme et al. (2016). In the forthcoming Sect. 3.5, we iteratively refit the trap release times, using these values (Table 4) as initial guesses.
Table 3: First estimates of the decay and release time constants for the identified trap species, obtained during Step 1. The species are ordered from fastest to slowest. Uncertainties on the release time constants are given at the $1\sigma$ level. The qualitative classification is based on the characteristic release timescale. The decay constant $b_{k}$ is obtained from an exponential fit to the overscan rows and is related to the species-specific release time $\tau_{\mathrm{r},k}$ via Eq. 10.
@@
1 Fast $2.87\times 10^{1}$ $3.10\times 10^{-5}$ $\pm 4.29\times 10^{4}$
2 Interm. $4.39\times 10^{-1}$ $2.03\times 10^{-3}$ $\pm 9.80\times 10^{-5}$
3 Interm. $3.37\times 10^{-2}$ $2.64\times 10^{-2}$ $\pm 3.21\times 10^{-6}$
4 Slow $5.68\times 10^{-3}$ $1.57\times 10^{-1}$ $\pm 9.98\times 10^{-9}$
Figure 8: First estimates of the release time constants obtained by fitting the overscan signal. The upper panel shows the measured signal plotted against the exponential model (Eq. 9) as a function of row index. The lower panel displays the residuals, indicating good agreement between the data and the model.
@@ 3.4 Smearing estimation and removal (Step 2)
Figure 9: Residuals from the smearing fit described in Sect. 3.4, shown for representative columns as a function of overscan row index.
@@ 3.5 Trap density calibration and radial dependence (Step 3)
The iterative strategy described in Sect. 3.3 determines the number of species to consider and their corresponding release times. We proceed to the next stage of calibration, which involves estimating the cumulative trap density via polynomial coefficients, after smearing removal (see previous section).
This stage used the Extended Pixel Edge Response (EPER) method (Prod’homme et al., 2016) for PLATO to analyse the overscan region. EPER relies on the CDM: the deferred charge appearing as trails in the overscan is modelled and fitted to the signal in the overscan region in order to infer the CTI parameters. From this, we derived the cumulative trap density for each species, $\mathrm{cnt}_{t,k}$, and constructed a model based on the best-fit values.
Based on the axisymmetric radial dependence inferred from the work of OHB System AG (2024) and covered in Sect 2.4, we modelled the mean trap density as a quadratic polynomial function
$\tilde{n}_{t,k}[i,j]=\sum_{p=0}^{N}a_{p,k}\,R_{i,j}^{\,p}$ (12)
where $p$ denotes the degree of the polynomial, $R$ denotes the normalised radial distance from the optical centre (see Eq. 2) and $i,j$ represent the row and column indices of the CCD, respectively.
From Eq. 12, we can similarly model the cumulative trap density as
@@
Eq. 13 expresses the cumulative number of traps $\tilde{\mathrm{cnt}}_{k}[i,j]$ for species $k$ at any given pixel $(i,j)$ as the sum of the local trap densities $\tilde{n}_{t,k}[i^{\prime},j]$ integrated along the parallel transfer direction up to row $i$. The specific case of $i=4510$ corresponds to the total number of CCD rows, representing full-column integration across the detector height. The cumulative number of traps is directly related to the radial dependence expressed by the polynomial coefficients $a_{p,k}$ from Eq. 12.
We implemented an iterative calibration algorithm applied simultaneously to the CCD image. The impact of CTI, and consequently the charge trails in the overscan rows, is primarily caused by the capture of charges in pixels located within a limited distance from the top of the CCD. Therefore, it is not necessary to model all the CCD rows, as doing so would significantly increase the computational cost without improving accuracy. To optimise the calibration process, we determined the number of science rows that must be considered by analysing the characteristic length $\mathrm{CL}$ of the effect of the slowest trap species. This characteristic length $\mathrm{CL}$ (in pixels) is defined as
$\mathrm{CL}=\frac{1}{b_{k}^{\mathrm{slowest}}}$ (14)
@@
For the CCD image part spanning the sum of $\mathrm{CL}_{\mathrm{eff}}$ and the overscan region, the Massey et al. (2014) correction is first applied to estimate the pixel intensities in the absence of CTI. This correction is used solely to initialize the estimation of the polynomial coefficients used to model the cumulative trap density, $\mathrm{cnt}_{k}$, as part of the calibration process. The procedure uses the run_cdm_cumul_radia_poly routine (see a Python implementation in Appendix A), which requires the current estimates of the polynomial coefficients $a_{p,k}$ for the working region, along with all the other constant CTI parameters.
Using the corrected pixel values over the full working region (i.e. $\mathrm{CL}_{\mathrm{eff}}+$ overscan), we forward-modelled CTI trailing to generate a synthetic “model” image, including the overscan. A cost function is then defined over the overscan region to optimize the polynomial coefficient map $a_{p,k}(i,j)$, the release time constants $\tau_{\mathrm{r,k}}$ and, the well-fill power-law index $\beta$, by minimizing the difference between the measured deferred charge and that predicted by the CDM (EPER method).
The release time associated with the first species has been fixed to $\tau_{\mathrm{r,0}}=3.05\times 10^{-5}\,\mathrm{s}$, corresponding to the value obtained in Step 1 (Table 3). This choice is motivated by the fact that $\tau_{\mathrm{r,0}}$ is shorter than the line transfer time ($\sim 900\penalty 10000\ \mu\mathrm{s}$), and therefore cannot be constrained by the fit.
The cost function and the CTI correction algorithm were combined in an iterative process: at each step, the corrected working region is updated, and the least-squares fit is repeated until the improvement in the reduced $\chi^{2}$ falls below $0.001$. To accelerate computations, we parallelized this process and selected every $N_{c}$ column within the working region to reduce the computational load.
@@
There is a total of 12 polynomial coefficients: three (for a quadratic function) per species across four trap species. The fit was initialized using the best-fit release times from Table 3, and a uniform starting guess of $1.0$ was adopted for each coefficient of the radial polynomial associated with the four trap species. We specified an upper bound of $\infty$ for all free parameters (except $\beta$). The lower bounds for coefficients $a_{0}$ and $a_{2}$ were set to zero, as these terms are not expected to be negative. However, the lower bound for the linear term $a_{1}$ was set to -$\infty$, motivated by the presence of a local maximum at the centre of the focal plane in Fig. 13, followed by a decrease and then a rise toward the edges. This pattern justified allowing $a_{1}<0$.
For the well-fill power index $\beta$, we used an initial guess of 0.5—representing the midpoint of its expected range between 0 and 1—and constrained it within the optimizer using bounds of 0 (lower) and 1 (upper). The initial guess for each release time $\tau_{\mathrm{r},k}$ corresponded to the values derived during Step 1 (see Table 3), with bounds of 0 and $\infty$ applied for the fitting. This iterative scheme yielded a map of the species-specific polynomial coefficients $a_{p,k}$, as well as fitted values for the release times $\tau_{\mathrm{r},k}$ and $\beta$, along with associated $1\sigma$ uncertainty maps. The final fit achieved a reduced $\chi^{2}$ of $1.223$. Figure 10 shows an example fit for a representative column (index 2255), located near the centre of the CCD.
We report the fitted values of the polynomial coefficients $a_{p,k}$, the release time constants $\tau_{\mathrm{r},k}$, and $\beta$ in Table 4, along with their respective $1\sigma$ uncertainties. The fitted value of $\beta$ is found to be very close to the true value from Prod’homme et al. (2016). The release time constants $\tau_{\mathrm{r},k}$ for species 3 and 4 are also estimated more accurately here than in Step 1, as shown in Table 4.
Figure 10: Iterative EPER-based fit for $\beta$, release time constants $\tau_{\mathrm{r,k}}$and polynomial coefficients $a_{\mathrm{p,k}}$ representing the trap densities in the overscan region for column index $2255$, with residuals shown in the lower panel.

### v2

@@ 2.5 Bias on transit depth measurements
To assess the impact of CTI on the transit method, we simulated representative light curves for both Earth- and Jupiter-sized planets transiting a solar-type star. The intrinsic stellar transit signal is generated using the PLATO Solar-like Light-curve Simulator [^3] https://sites.lesia.obspm.fr/psls/ (Samadi et al., 2019), which is based on the analytical model by Mandel and Agol (2002) and its Python implementation by Ian Crossfield (MIT) [^4] http://www.astro.ucla.edu/~ianc/.
The flux variations over time are interpolated onto a grid of 2000 time steps. The stellar image is modelled using a Gaussian point spread function (PSF) with a characteristic width of $0.6$ pixel. The total flux is scaled to match the expected value for a given stellar magnitude, following the flux–magnitude relation from Marchiori et al. (2019). Finally, this flux distribution is modulated by the intrinsic transit light curve.
We added a constant background contribution from zodiacal light of $45\penalty 10000\ \mathrm{e^{-}\,s^{-1}\,pixel^{-1}}$ (Marchiori et al., 2019). Given the integration time of $21$s for the PLATO CCD, the net background value is $\approx 1000\penalty 10000\ \mathrm{e^{-}\,pixel^{-1}}$. A higher background fills traps that would otherwise capture flux from the target star, thereby reducing the CTI effect. The aperture size is determined using our optimal aperture mask routine, following the approach of Marchiori et al. (2019). Finally, from the simulated light curve, a transit depth is extracted.
@@
We calculated the bias on transit depth due to CTI for cases corresponding to nominal mission lifetime of 4 years and extended mission lifetime of 8 years. We also incorporated positional dependence, based on the imagette’s location in a region of high or low CTI. We evaluated three representative CCD positions, bottom left (A), top left (B), and top right (C), corresponding to different trap density regimes derived from Eq.7 (see Fig.4). The bottom-left corner (A) is the region with the lowest cumulative trap density. Conversely, the top-right corner (C) has the highest cumulative trap density. The top-left corner (B) represents an intermediate case.
@@
To quantify the impact of CTI on transit light curves, we defined the photometric bias, $\Delta$, as the difference between the pristine transit depth (no CTI) and the CTI-affected depth, expressed in both pixels-per-million (ppm) and percentage. Table 2 reports the impact of CTI on the transit depth of an Earth-sized planet transiting a solar-type star with $m_{V}=11$ under different mission durations and trap density scenarios. For completeness, it also presents the corresponding bias for a bright star ($m_{V}=9$) and a faint star ($m_{V}=13$) in a region of high CTI (position C) during the extended 8-year mission. It is worth noting that these calculations were done for the worst CCD (among all the cameras), i.e. for the CCD with the highest maximum in the TNID.
Table 2: CTI depth across mission durations, detector positions, and magnitudes. [^5] The pristine transit depth is equal to $105.8$ ppm (typical value for an Earth-sized planet orbiting a solar-type star).
@@
The origin of this bias is the CTI-induced charge loss, which does not scale linearly with the stellar flux. As a result, the deferred charge release affects the in-transit and out-of-transit measurements differently, thereby influencing the measured transit depth. This non-linear behaviour is the root cause of the photometric bias seen in the observed light curve.
We found that the horizontal drift had no impact on the bias, as the light curves had been de-trended beforehand. This will also be the case for the PLATO light curves (however, see the discussion in Sect. 5.2).
Figure 5: Comparison of transit light curves for the worst-case scenario: an 8-year mission at high trap density (position C) and $m_{V}=11$. The solid orange curve includes CTI effects, while the dashed blue curve shows the CTI-free case. CTI-induced charge loss introduces a photometric bias of $3.95\%$.
PLATO science requirements dictate that the planet-to-star radius ratio ($R_{\mathrm{p}}/R_{\mathrm{s}}$) must be measured with a relative precision better than $2\%$ for an Earth-sized planet transiting a G-type star (goal: $m_{V}=11$; PLATO Science Working Team 2019). Because the transit depth ($\delta$) is proportional to the square of this ratio, standard error propagation yields
@@ 3 CTI calibration strategy
This section outlines the calibration strategy employed in this study, which relies on full or partial images acquired at the beginning of each quarter (i.e. every three months). First, we defined an empirical model to determine the number of trap species and to derive initial estimates for their release time constants ($\tau_{\mathrm{r,k}}$). Next, we addressed and removed the contribution from smearing. Finally, we adopted a polynomial model to describe the spatial variation in the trap density, reproducing the radial dependence discussed in Sect. 2.4. We then applied the EPER method (Prod’homme et al., 2016) to constrain the polynomial coefficients of the trap density model ($a_{\mathrm{p,k}}$), along with the $\beta$ parameter and the release time constants ($\tau_{\mathrm{r,k}}$). The overall procedure is summarised in Fig. 6.
Figure 6: Workflow of the CTI calibration strategy: initial estimates for the release time constants ($\tau_{\mathrm{r,k}}$) are obtained by fitting an empirical model to the overscan signal; smearing is corrected via a column-wise exponential-plus-constant fit; the mean trap density is modelled as a quadratic polynomial to reproduce its radial dependence; and the EPER method is applied, in conjunction with a CTI correction algorithm, to iteratively fit the polynomial coefficients and the well-fill power-law index ($\beta$), and to refine the release time constants ($\tau_{\mathrm{r,k}}$).
@@ 3.1 Calibration dataset
As PLATO is not yet in-flight, we used a simulated CCD image containing a stellar field for the calibration procedure. This image is generated with PLATOSim (Jannsen et al., 2024), a simulation tool designed to produce photometric time series and CCD images tailored to the PLATO mission. The simulated image includes several sources of noise—smearing, photon noise, and readout noise—to approximate realistic observational conditions. To improve the signal-to-noise ratio (S/N), we averaged over 10 exposures. The upper limit of 10 exposures is decided by the in-flight operational constraints (specifically to minimise the duration of the calibration phase and the associated telemetry volume).
@@ 3.2 Overscan analysis and CTI signature
Parallel overscan rows are virtual rows beyond the CCD’s physical imaging area, accessed after the last physical row is read out. Hereafter, ‘overscan’ refers to this parallel overscan region. These rows contain deferred charges that were not recorded during readout because of trap capture, providing a diagnostic window to constrain the CTI signature. Figure 7 shows the overscan signal, summed across all the columns for a given row index. Its shape reflects the delayed release of charge by different trap species: those with short release times dominate near the start, whereas slower species extend further into the overscan region. The slowest traps are of primary concern, as their release occurs beyond the photometric aperture.
Figure 7: Sum of the signal in each overscan row across all columns as a function of row index. The profile follows an exponential decay, with physical rows located before overscan row index 0.
@@ 3.3 Initial estimates for the trap release time (Step 1)
From the profile of the signal in Fig. 7, a multi-exponential function is used to (empirically) model the signal contained in the overscan region. It is illustrated in the Eq. 9 as
@@
where $A_{k}$ is the amplitude for trap species $k$, corresponding to the range (max - min) of the signal; $b_{k}$ is the decay rate for species $k$, inversely related to the release time (see Eq. 10); $C$ is the constant offset (sum of smearing in each individual column); $i$ is the row index along the CCD column (parallel transfer direction); and $n$ is the total number of trap species; $Y_{\text{CTI}}(i)$ is the signal level (in e-/pix) at row $i$, including CTI trailing effects. The relationship between the decay rate ($b_{k}$) and release time (${\tau}_{r,k}$) is given by
${\tau}_{r,k}=\frac{1}{b_{k}}\times t_{\text{transfer\ cadence}}\,.$ (10)
We designed an iterative fitting strategy that progressively increases the number of species until the solution converges within a predefined threshold. Applying this strategy to our simulated image yielded a maximum of four trap species, with a final reduced $\chi^{2}$ of $1.141$. The first estimates for the release time values and decay constants for these species, ordered from fastest to slowest, are listed in Table 4. We provide the associated uncertainties for the release times and classify each species according to its characteristic release timescale: fast, intermediate, or slow. The corresponding fit results, along with the residuals, are presented in Fig. 8. The difference between our best-fit release time values (Table 4) and the reference values from Prod’homme et al. (2016, see our Table 1) arises from the distinct methodologies employed in each case. Our simulations include a realistic stellar field, which influences the empirical model used to derive best-fit release times, unlike the flat-field image adopted by Prod’homme et al. (2016). In the forthcoming Sect. 3.5, we iteratively refit the trap release times, using these values (Table 4) as initial guesses.
Table 3: First estimates of decay and release time constants obtained during Step 1.666The species are ordered from fastest to slowest. Uncertainties on the release time constants are given at the $1\sigma$ level. The qualitative classification is based on the characteristic release timescale. The decay constant ($b_{k}$) is obtained from an exponential fit to the overscan rows and is related to the species-specific release time ($\tau_{\mathrm{r},k}$) via Eq. 10. Uncertainty for Species 1 is marked with a dash as it cannot be constrained because its release time is shorter than the line transfer time.
@@
1 Fast $2.87\times 10^{1}$ $3.10\times 10^{-5}$ -
2 Interm. $4.39\times 10^{-1}$ $2.03\times 10^{-3}$ $\pm 4.26\times 10^{-5}$
3 Interm. $3.37\times 10^{-2}$ $2.64\times 10^{-2}$ $\pm 5.44\times 10^{-4}$
4 Slow $5.68\times 10^{-3}$ $1.57\times 10^{-1}$ $\pm 3.44\times 10^{-3}$
Figure 8: First estimates of the release time constants obtained by fitting the overscan signal. Top: Measured signal plotted against the exponential model (Eq. 9) as a function of row index. Bottom: Residuals, indicating good agreement between the data and the model.
@@ 3.4 Smearing estimation and removal (Step 2)
Figure 9: Residuals from the smearing fit described in Sect. 3.4, shown as a function of column index.
@@ 3.5 Trap density calibration and radial dependence (Step 3)
The iterative strategy described in Sect. 3.3 determines the number of species to consider and their corresponding release times. The next stage of calibration involved estimating the cumulative trap density via polynomial coefficients, after smearing removal (see the previous section).
This stage used the EPER method (Prod’homme et al., 2016) for PLATO to analyse the overscan region. EPER relies on the CDM: the deferred charge appearing as trails in the overscan is modelled and fitted to the signal in the overscan region in order to infer the CTI parameters. From this, we derived the cumulative trap density for each species, $\mathrm{cnt}_{t,k}$, and constructed a model based on the best-fit values.
Based on the axisymmetric radial dependence inferred from the work of OHB System AG (2024) and covered in Sect. 2.4, we modelled the mean trap density as a quadratic polynomial function:
$\tilde{n}_{t,k}[i,j]=\sum_{p=0}^{N}a_{p,k}\,R_{i,j}^{\,p}\,$, (12)
where $p$ denotes the degree of the polynomial, $R$ denotes the normalised radial distance from the optical centre (see Eq. 2), and $i$ and $j$ represent the row and column indices of the CCD, respectively. From Eq. 12, we can similarly model the cumulative trap density as
@@
Equation 13 expresses the cumulative number of traps ($\tilde{\mathrm{cnt}}_{k}[i,j]$) for species $k$ at any given pixel $(i,j)$ as the sum of the local trap densities ($\tilde{n}_{t,k}[i^{\prime},j]$) integrated along the parallel transfer direction up to row $i$. The specific case of $i=4510$ corresponds to the total number of CCD rows, representing full-column integration across the detector height. The cumulative number of traps is directly related to the radial dependence expressed by the polynomial coefficients ($a_{p,k}$) from Eq. 12.
We implemented an iterative calibration algorithm applied simultaneously to the CCD image. The impact of CTI, and consequently the charge trails in the overscan rows, is primarily caused by the capture of charges in pixels located within a limited distance from the top of the CCD. Therefore, it is not necessary to model all the CCD rows, as doing so would significantly increase the computational cost without improving accuracy. To optimise the calibration process, we determined the number of science rows that must be considered by analysing the characteristic length ($\mathrm{CL}$) of the effect of the slowest trap species. This characteristic length $\mathrm{CL}$ (in pixels) is defined as
$\mathrm{CL}=\frac{1}{b_{k}^{\mathrm{slowest}}}\,$, (14)
@@
For the CCD image part spanning the sum of $\mathrm{CL}_{\mathrm{eff}}$ and the overscan region, the Massey et al. (2014) correction is first applied to estimate the pixel intensities in the absence of CTI. This correction is used solely to initialise the estimation of the polynomial coefficients used to model the cumulative trap density, $\mathrm{cnt}_{k}$, as part of the calibration process. The procedure uses the run_cdm_cumul_radia_poly routine (see the Python implementation in plato-cti-correction), which requires the current estimates of the polynomial coefficients ($a_{p,k}$) for the working region, along with all the other constant CTI parameters.
Using the corrected pixel values over the full working region (i.e. $\mathrm{CL}_{\mathrm{eff}}+$ overscan), we forward-modelled CTI trailing to generate a synthetic ‘model’ image, including the overscan. A cost function is then defined over the overscan region to optimise the polynomial coefficient map ($a_{p,k}(i,j)$), the release time constants ($\tau_{\mathrm{r,k}}$), and the well-fill power-law index ($\beta$), by minimising the difference between the measured deferred charge and that predicted by the CDM (EPER method).
The release time associated with the first species has been fixed to $\tau_{\mathrm{r,0}}=3.10\times 10^{-5}\,\mathrm{s}$, corresponding to the value obtained in Step 1 (Table 3). This choice is motivated by the fact that $\tau_{\mathrm{r,0}}$ is shorter than the line transfer time ($\sim 900\penalty 10000\ \mu\mathrm{s}$), and therefore cannot be constrained by the fit.
The cost function and the CTI correction algorithm were combined in an iterative process: at each step, the corrected working region is updated, and the least-squares fit is repeated until the improvement in the reduced $\chi^{2}$ falls below $0.001$. To accelerate computations, we parallelised this process and selected every $N_{c}$ column within the working region to reduce the computational load.
@@
There is a total of 12 polynomial coefficients: three (for a quadratic function) per species across four trap species. The fit was initialised using the best-fit release times from Table 3, and a uniform starting guess of $1.0$ was adopted for each coefficient of the radial polynomial associated with the four trap species. We specified an upper bound of $\infty$ for all free parameters (except $\beta$). The lower bounds for coefficients $a_{0}$ and $a_{2}$ were set to zero, as these terms are not expected to be negative. However, the lower bound for the linear term $a_{1}$ was set to -$\infty$, motivated by the presence of a local maximum at the centre of the focal plane in Fig. 13, followed by a decrease and then a rise towards the edges. This pattern justified allowing $a_{1}<0$.
For the well-fill power index ($\beta$), we used an initial guess of 0.5—representing the midpoint of its expected range between 0 and 1—and constrained it within the optimiser using bounds of 0 (lower) and 1 (upper). The initial guess for each release time ($\tau_{\mathrm{r},k}$) corresponded to the values derived during Step 1 (see Table 3), with bounds of 0 and $\infty$ applied for the fitting. This iterative scheme yielded a map of the species-specific polynomial coefficients ($a_{p,k}$), as well as fitted values for the release times ($\tau_{\mathrm{r},k}$) and $\beta$, along with associated $1\sigma$ uncertainty maps. The final fit achieved a reduced $\chi^{2}$ of $1.223$. Figure 10 shows an example fit for a representative column (index 2255), located near the centre of the CCD.
We report the fitted values of the polynomial coefficients ($a_{p,k}$), the release time constants ($\tau_{\mathrm{r},k}$), and $\beta$ in Table 4, along with their respective $1\sigma$ uncertainties. The fitted value of $\beta$ is found to be very close to the true value from Prod’homme et al. (2016). The release time constants ($\tau_{\mathrm{r},k}$) for species 3 and 4 are also estimated more accurately here than in Step 1, as shown in Table 4.
Figure 10: Top: Iterative EPER-based fit for $\beta$, release time constants ($\tau_{\mathrm{r,k}}$), and polynomial coefficients ($a_{\mathrm{p,k}}$) representing the trap densities in the overscan region for column index $2255$. Bottom: Residuals.